### v3

@@ Abstract
In this work, we propose an adjoint-based optimization procedure to control the onset of the Rayleigh-Bénard instability with a melting front. A novel cut cell method is used to solve the Navier-Stokes equations in the Boussinesq approximation and the convection-diffusion equation in the fluid layer, as well as the heat equation in the solid phase. To track the interface we use the level set method where its evolution is simply governed by an advection equation. An incomplete continuous adjoint problem is then derived by considering that the velocity field is a check-pointing variable. The results of the minimization problem with a tracking-type cost-functional show that our adjoint method is well suited to optimize the shapes of the fronts in this configuration.
@@ 1 Introduction
Stefan problem [1] models transport and transfer phenomena, in particular solid-liquid phase change in evaporating or chemically reacting flows. Such phenomena govern the interface motion in many engineering related problems such as dendritic solidification [2, 3], phase transformation in metallic alloys [4], and solid fuel combustion [5]. At the interface, the Stefan condition arises from the interaction of liquid and solid phases (both considered incompressible), resulting in a moving liquid-solid interface (freezing or melting front). The speed of the front is directly related to the jump in the conductive heat flux across the interface. In one dimension, this problem has been studied in depth [6, 7, 8]; in higher dimensions, however, due to the unstable pattern formation [9, 10, 11] specific numerical methods, such as the level set method [12, 2, 13, 14, 15, 16, 17, 18, 19, 20, 21], have been used to predict the motion and shape of the interface. In crystal growth, for example, under-cooling triggers an instability mechanism, causing the solid phase of the material to grow into the liquid phase in a finger-like or dendritic fashion, resulting in complex interfacial shapes and possible topography changes. Moreover, while convection with a fixed topography is usually studied in the classical Rayleigh-Bénard setup [22], many natural phenomena such as erosion or melting involve a coupling between the flow and the moving boundary [23, 24].
@@
The shape of the resulting interface may strongly affect the outcome and time-frame of the production processes in many industrial applications involving phase change [25, 26]. As a result, it is desirable to extract efficient control strategies to manipulate the motion of the interface, for instance, by tracking a prescribed trajectory. Two major types of optimization methods in use today are (i) gradient-based, and (ii) derivative free methods. While an efficient class of generic algorithms (belonging to the class of derivative free methods) based on the surrogate management framework [27] and artificial neural networks [28] have been used for optimization in fluid mechanics, mainly in the area of aerodynamic shape optimization, they could require many function evaluations, for training purposes for example. When detailed simulations of interfacial flows are concerned, each function evaluation commands a full (potentially unsteady) CFD computation, causing gradient-based methods to be at an advantage. Most common methods in extracting the gradient information, on the other hand, are analytical or use finite differences, neither of which are suited to the configuration of interest to this work. Adjoint-based algorithms present a suitable alternative, as they allow the determination of the gradient at a cost comparable to a single function evaluation [29, 30, 31, 32]. The use of adjoint methods for design and optimization has been an active area of research which started with the pioneering work of Pironneau [33] with applications in fluid mechanics, and has been extensively used in aeronautical shape optimization by Jameson and co-workers [34, 35]. Ever since these groundbreaking studies, adjoint-based methods have been widely used in fluid mechanics particularly in the areas of aero- and thermo-acoustics [36, 37]. More recently flow regimes dominated by nonlinear dynamics, such as separation and mixing have also been analysed using adjoint-based techniques [38, 39, 40, 41]. Adjoint-based methods have also been employed for the purpose of sensitivity analysis or control in flows in the presence of large gradients (flames or interfaces) [42, 43, 44, 45], showing great promise, and therefore are adopted here to carry out the optimization procedure.
@@
In the context of Stefan problems various control strategies have been employed to track the location of the interface. In a one-dimensional setting, for example, set-valued fixed point equations [46] or linear-quadratic defect minimization [47] have been used to control the location of the front. Adjoint-based algorithms have also been applied to a Stefan problem caused by heterogeneous reactions on a surface of a one-dimensional solid particle [5] to extract sensitives with respect to various kinetic parameters. Alternatively, in a two-dimensional setting, adjoint-based algorithms have been utilized previously together with finite element and finite difference approaches to track and control the location of interface by imposing heat flux (or temperature) at the boundary in order to realize the desired interface motion [48, 49, 50]. In particular, Bernauer & Herzog [51], making use of shape calculus tools, derived the set of adjoint equations to extract control strategies for Stefan problems with a sharp representation of the interface.
@@
Similar to the approach of [51] shape calculus tools have been employed to extract the corresponding adjoint equations in our previous work [17]. However, contrary to the previous studies, control strategies are extracted here to control the interface shape resulting from the melting boundary. In this case, an incomplete adjoint – by considering the velocity field in the fluid layer as a check-pointing variable – is derived and used to minimize a tracking-type cost functional through a gradient-based algorithm. The velocity field is treated as a known variable in the adjoint convection-diffusion equation. We will show that the derived incomplete adjoint is sufficient to retrieve meaningful information of the system and drive the cost functional towards its minimum.
@@ 2 Governing equations
Following the recent studies on the Rayleigh-Bénard instability in the presence of a melting boundary [52, 12], we consider the evolution of an initially flat liquid layer heated from below and comprised between a wall and a solid layer. As the solid melts, the liquid layer will grow vertically until a critical height where convection kicks in. A schematic of this configuration is shown in Fig. 1. The gravity is pointing downwards $\mathbf{g}=-g\mathbf{e}_{y}$ and the horizontal size of the domain is $bH$, where $b$ is the aspect ratio and where $H=1$ is the height of the system. The temperature at the bottom wall is $T_{b}=0.7$, and the one at the top wall is our control variable $w(x)<0$, such that $w(x)<T_{M}<T_{b}$ where $T_{M}=0$ is the melting temperature at the front. The choice of $T_{b}$ is such that $\Delta T=1$ when w(x)=0.3w(x)=-0.3. The initial flat interface is located at $h_{0}=0.05$ and the initial temperature field is set to zero.
@@
Finally, as phase change occurs, there will appear a latent heat which is either absorbed or released. The condition of heat conservation at a given point on the moving interface corresponds to the rate at which heat is generated at the boundary, balanced by the rate at which this heat flows in either phase. Along the interface, the Stefan condition [53] states that
@@
where $L_{H}$ is the latent heat. The moving interface is captured using the level set method. A level set function $\phi$ [54] is defined on the computational domain $\Omega$ to map the locus of one of its iso-levels ($\left\{\left.\mathbf{x}\in\Omega\right|\phi\left(\mathbf{x},t\right)=\phi_{0}\right\}$) to an interface $\Gamma\left(t\right)$ that separates two non-overlapping domains, $\Omega_{S}\left(t\right)$ and $\Omega_{L}\left(t\right)$, each occupied by a different phase. The value $\phi$ is defined as the signed distance to the interface,
@@
provided that the velocity field can be smoothly extended from $\Gamma\left(t\right)$. The most natural algorithm is to let $v$ be a constant along the lines normal to $\Gamma$. To achieve this, the method described in [55] is adopted here. Using this approach, the velocity is extended in the normal direction by solving the following hyperbolic partial differential equation
@@
where $\Delta$ is the uniform grid spacing, $(x)^{+}=\operatorname{max}(0,x)$ and $(x)^{-}=\operatorname{min}(0,x)$, and the time step $\tau^{\star}$ is chosen so that $\tau^{\star}/\Delta^{2}=0.45$. The pseudo-time spawn in the velocity extension algorithm is purely fictitious and the number of iterations in Equation 14 corresponds to the width of the narrow-band (NB) around the 0-level set where the velocities are initialized. A semi-implicit scheme described in [56, 57] is used here to solve the level set advection equation. Using this method, the equation is written in an alternative form
@@
In order to discretize Equation 16, we use the same notation as in [56]. Consider $\mathrm{p}$ to be a finite volume of a cell and $e_{\mathrm{p}\mathrm{q}}$ the edge between $\mathrm{p}$ and $\mathrm{q}$, $\mathrm{q}\in\mathcal{N}(\mathrm{p})$ where $\mathcal{N}(\mathrm{p})$ is the set of neighboring finite volumes. The length of $e_{\mathrm{p}\mathrm{q}}$ is normalized to $1$. Let $\mathbf{n}_{\mathrm{p}\mathrm{q}}$ be the outer normal vector to $e_{\mathrm{p}\mathrm{q}}$ with respect to $\mathrm{p}$. Finally, let us denote $\bar{\phi}_{\mathrm{p}}$ the constant reconstruction of $\phi_{\mathrm{p}}$ in the finite volume $p$ and $\bar{\phi}_{\mathrm{p}\mathrm{q}}$ the constant reconstruction of $\phi$ on $e_{\mathrm{p}\mathrm{q}}$. Integrating volume $\mathrm{p}$, Equation 16 yields
@@
where $\tau$ is the time step, $\Delta$ the uniform grid spacing and $n$ a given time step. To reconstruct of $\bar{\phi}_{\mathrm{p}}$, $\bar{\phi}_{\mathrm{p}\mathrm{q}}$ and $|\nabla\phi_{\mathrm{p}\mathrm{q}}|$, we use the diamond-cell strategy described in [57]. This evolution equation however does not preserve the signed distance property, which can be restored by periodically employing a reinitialization procedure by solving an Eikonal equation. There exist many numerical methods for solving this equation and, here we have adopted that of Min [58], also recently used in the work of Limare [12]. For further details on the numerical implementation of the level set coupled to a cut cell method, we refer the reader to [17, 19, 20].
@@ 3 Adjoint problem
We will now proceed with the derivation of the corresponding continuous adjoint problem. The main novelty behind our derivation is that the velocity field $\mathbf{u}$, known at each time step, can be considered as a check-pointing variable, therefore discarding the adjoint of the of the Navier-Stokes equations. Indeed, as phase-change is the main driver of this phenomena, one can neglect the velocity in the adjoint. The minimization procedure will serve as a validation of this hypothesis.
@@
Turning now to the adjoint level set problem; due to the geometric non-linearity induced by $\phi$ on the domains of integration a careful treatment of each term is required as described in [51]. Denoting $D\llbracket H(\phi);\delta\phi\rrbracket$ the variation of $H(\cdot)$ in the direction $\delta\phi$, the adjoint level set equations are derived by setting $\mathcal{L}_{\phi}(\cdot)=0$
@@
This equation can be interpreted as a first-order conservation law on $\Gamma$ and the source term on the right-hand-side can be extended in the neighborhood of $\Gamma$. For details on the derivation and simplifications of $\mathcal{L}_{\phi}(\cdot)=0$, we refer the reader to [17].
@@ 4 Optimization results
Adjoint-based methods—being efficient gradient-based optimization approaches—are often the methods of choice in PDE-constrained optimal control. By solving an adjoint equation once, one obtains the gradient of the objective with respect to all control variables at minimal cost, making the optimal-control formulation both tractable and efficient even for large-scale parameter spaces. We choose to solve the minimization problem (MP) by using the limited memory BFGS (L-BFGS) method, a quasi-Newton method originally described in [59]. The main characteristic of this method is that it determines the descent direction by preconditioning the gradient with an approximation of the Hessian matrix. This information is obtained using past approximations – the number of approximations is determined by the memory length parameter which is set to $m=10$ – as well as the gradient. As an initial guess for the initial Hessian, we use the scaled identity matrix as described in [60]. Algorithm 1 summarizes the L-BFGS method used in our numerical example. The algorithm is stopped at a given iteration $n$ if one of the following criteria is fulfilled
@@
   Solve the forward Stefan problem (2) for $T^{k}$ and $\phi^{k}$
   Solve the adjoint Stefan problem (3) for $\Theta^{k}$ and $\psi^{k}$
   Compute the gradient:
@@
  Choose an initial approximation to the inverse of the Hessian $H^{k}_{0}$ Construct the direction $d^{k}=-H^{k}\nabla\mathcal{J}^{k}$ Determine $\sigma^{k}$ using a Line Search algorithm with backtracking where $\sigma^{k}=\operatorname{argmin}\mathcal{J}(w^{k}+\sigma^{k}d^{k})$ Update $w^{k+1}=w^{k}+\sigma^{k}d^{k}$ $k\rightarrow k+1$
@@
We now compare the results to a derivative-free method (where no information on the gradient is required), the Particle Swarm optimization algorithm [61]. This method attempts to improve global convergence by switching between four evolutionary states: exploration, exploitation, convergence, and jumping out. In the jumping out state it tries to take the best particle and move it away from its local optimum, to improve the ability to find a global one. We test the derivative-free optimization procedure on the cases presented using the same parameters for the cost functional. Table 1 summarizes the comparison between both methods. As expected, the derivative-free one attains a lower minimum at a cost of a much higher number of function evaluation (one order of magnitude higher).

### v4

@@ Abstract
In this work, we propose an adjoint-based optimization procedure to control the onset of the Rayleigh-Bénard instability with a melting front. A novel cut cell method is used to solve the Navier-Stokes equations in the Boussinesq approximation and the convection-diffusion equation in the fluid layer, as well as the heat equation in the solid phase. To track the interface we use the level set method where its evolution is simply governed by an advection equation. An incomplete continuous adjoint problem is then derived by treating the velocity field obtained from the forward problem as a known variable in the adjoint convection-diffusion equation, thereby avoiding the need to solve a Navier–Stokes adjoint in the fluid phase. To the best of our knowledge, this provides the first adjoint-based optimization framework for Rayleigh–Bénard instability with a melting boundary. Two optimization problems, together with a comparison against a derivative-free particle–swarm method, demonstrate that the proposed incomplete adjoint yields gradients accurate enough to control the front shape while reducing the number of expensive function evaluations by about an order of magnitude.
@@ 1 Introduction
Stefan problem [sarler_stefans_1995] models transport and transfer phenomena, in particular solid-liquid phase change in evaporating or chemically reacting flows. Such phenomena govern the interface motion in many engineering related problems such as dendritic solidification [osher_fronts_1988, Juric1996], phase transformation in metallic alloys [Segal1998], and solid fuel combustion [Hassan2021]. At the interface, the Stefan condition arises from the interaction of liquid and solid phases (both considered incompressible), resulting in a moving liquid-solid interface (freezing or melting front). The speed of the front is directly related to the jump in the conductive heat flux across the interface. In one dimension, this problem has been studied in depth [javierre_comparison_2006, Rose1993, Brattkus1992]; in higher dimensions, however, due to the unstable pattern formation [langer_instabilities_1980, mullins_stability_1964, woods_melting_1992] specific numerical methods, such as the level set method [Limare2022, osher_fronts_1988, gibou_level_2003, chen_numerical_2009, jenkins_level_2015, bayat_sharp_2022, fullana2022, silva_topology_2022, Fullana2023, QuirosRodriguez2024, castanar_topology_2024], have been used to predict the motion and shape of the interface. In crystal growth, for example, under-cooling triggers an instability mechanism, causing the solid phase of the material to grow into the liquid phase in a finger-like or dendritic fashion, resulting in complex interfacial shapes and possible topography changes. Moreover, while convection with a fixed topography is usually studied in the classical Rayleigh-Bénard setup [bodenschatz_recent_2000], many natural phenomena such as erosion or melting involve a coupling between the flow and the moving boundary [Yang_Howland_Liu_Verzicco_Lohse_2023, perissutti2024morphodynamicsmeltingiceturbulent].
@@
The shape of the resulting interface may strongly affect the outcome and time-frame of the production processes in many industrial applications involving phase change [picelli_topology_2020, siqueira_topology_2024]. As a result, it is desirable to extract efficient control strategies to manipulate the motion of the interface, for instance, by tracking a prescribed trajectory. Two major types of optimization methods in use today are (i) gradient-based, and (ii) derivative free methods. While an efficient class of generic algorithms (belonging to the class of derivative free methods) based on the surrogate management framework [Marsden2008] and artificial neural networks [Pierret2007] have been used for optimization in fluid mechanics, mainly in the area of aerodynamic shape optimization, they could require many function evaluations, for training purposes for example. When detailed simulations of interfacial flows are concerned, each function evaluation commands a full (potentially unsteady) CFD computation, causing gradient-based methods to be at an advantage. Adjoint-based algorithms present a suitable method, as they allow the determination of the gradient at a cost comparable to a single function evaluation [Giles2000, griewank_algorithm_2000, rodrigues_adjoint-based_2020, burghardt_discrete_2022]. The use of adjoint methods for design and optimization has been an active area of research which started with the pioneering work of Pironneau [Pironneau1974] with applications in fluid mechanics, and has been extensively used in aeronautical shape optimization by Jameson and co-workers [Jameson88, Jameson98]. Ever since these groundbreaking studies, adjoint-based methods have been widely used in fluid mechanics particularly in the areas of aero- and thermo-acoustics [Juniper2010, Lemke2013]. More recently flow regimes dominated by nonlinear dynamics, such as separation and mixing have also been analysed using adjoint-based techniques [Schmidt2013, Rabin2014, Foures2014, Duraisamy2012]. Adjoint-based methods have also been employed for the purpose of sensitivity analysis or control in flows in the presence of large gradients (flames or interfaces) [Fikl2020, ou_unsteady_2011, Braman2015, Lemke2019], showing great promise, and therefore are adopted here to carry out the optimization procedure.
@@
In the context of Stefan problems various control strategies have been employed to track the location of the interface. In a one-dimensional setting, for example, set-valued fixed point equations [Hoffmann1982] or linear-quadratic defect minimization [Knabner1985] have been used to control the location of the front. Adjoint-based algorithms have also been applied to a Stefan problem caused by heterogeneous reactions on a surface of a one-dimensional solid particle [Hassan2021] to extract sensitives with respect to various kinetic parameters. Alternatively, in a two-dimensional setting, adjoint-based algorithms have been utilized previously together with finite element and finite difference approaches to track and control the location of interface by imposing heat flux (or temperature) at the boundary in order to realize the desired interface motion [Kang1995, Yang1997, Hinze2007]. In particular, Bernauer & Herzog [bernauer_optimal_2011], making use of shape calculus tools, derived the set of adjoint equations to extract control strategies for Stefan problems with a sharp representation of the interface.
@@
Similar to the approach of [bernauer_optimal_2011] shape calculus tools have been employed to extract the corresponding adjoint equations in our previous work [fullana2022]. However, contrary to the previous studies, control strategies are extracted here to control the interface shape resulting from the melting boundary. For the configuration considered here, deriving the fully coupled analytical adjoint of the Navier–Stokes–Stefan system on a moving domain is not tractable, which motivates the use of an incomplete continuous adjoint tailored to this problem. The incomplete adjoint is derived – by considering the velocity field in the fluid layer as a known variable in the adjoint convection-diffusion equation – and used to minimize a tracking-type cost functional through a gradient-based algorithm. We will show that the derived incomplete adjoint is sufficient to retrieve meaningful information on the system and to drive the cost functional towards its minimum at a substantially reduced computational cost.
The main contributions of this work can be summarized as follows. (i) A sharp-interface optimal-control problem for the Rayleigh–Bénard instability with a melting boundary, based on a level-set/cut-cell discretization of the two-phase Stefan problem under the Boussinesq approximation is formulated. (ii) We then derive an incomplete continuous adjoint that neglects the Navier–Stokes adjoint by treating the velocity field from the forward simulation as a known variable in the adjoint convection–diffusion equation. (iii) We demonstrate, on two non-trivial optimization problems and in comparison with a derivative-free method, that the resulting gradients are sufficiently accurate to control the front shape while significantly reducing the number of function evaluations.
The present study is restricted to a two-dimensional configuration with a sharp interface representation and without surface-tension effects. The control is exerted only through the temperature distribution prescribed at the upper boundary of a periodic domain. As such, the proposed framework is intended as a first step towards adjoint-based control of melting fronts in convection-driven problems.
@@ 2 Governing equations
Following the recent studies on the Rayleigh-Bénard instability in the presence of a melting boundary [Favier2019, Limare2022], we consider the evolution of an initially flat liquid layer heated from below and comprised between a wall and a solid layer. As the solid melts, the liquid layer will grow vertically until a critical height where convection kicks in. A schematic of this configuration is shown in Fig. 1. The gravity is pointing downwards $\mathbf{g}=-g\mathbf{e}_{y}$ and the horizontal size of the domain is $bH$, where $b$ is the aspect ratio and where $H=1$ is the height of the system. The temperature at the bottom wall is $T_{b}=0.7$, and the one at the top wall is our control variable $w(x)<0$, such that $w(x)<T_{M}<T_{b}$ where $T_{M}=0$ is the melting temperature at the front. The choice of $T_{b}$ is such that $\Delta T=1$ when w(x)=0.3w(x)=-0.3. The initial flat interface is located at $h_{0}=0.05$ and the initial temperature field is set to zero.
@@
Finally, as phase change occurs, there will appear a latent heat which is either absorbed or released. The condition of heat conservation at a given point on the moving interface corresponds to the rate at which heat is generated at the boundary, balanced by the rate at which this heat flows in either phase. Along the interface, the Stefan condition [Stefan1891] states that
@@
where $L_{H}$ is the latent heat. The moving interface is captured using the level set method. A level set function $\phi$ [Sethian1999] is defined on the computational domain $\Omega$ to map the locus of one of its iso-levels ($\left\{\left.\mathbf{x}\in\Omega\right|\phi\left(\mathbf{x},t\right)=\phi_{0}\right\}$) to an interface $\Gamma\left(t\right)$ that separates two non-overlapping domains, $\Omega_{S}\left(t\right)$ and $\Omega_{L}\left(t\right)$, each occupied by a different phase. The value $\phi$ is defined as the signed distance to the interface,
@@
provided that the velocity field can be smoothly extended from $\Gamma\left(t\right)$. The most natural algorithm is to let $v$ be a constant along the lines normal to $\Gamma$. To achieve this, the method described in [peng_pde-based_1999] is adopted here. Using this approach, the velocity is extended in the normal direction by solving the following hyperbolic partial differential equation
@@
where $\Delta$ is the uniform grid spacing, $(x)^{+}=\operatorname{max}(0,x)$ and $(x)^{-}=\operatorname{min}(0,x)$, and the time step $\tau^{\star}$ is chosen so that $\tau^{\star}/\Delta^{2}=0.45$. The pseudo-time spawn in the velocity extension algorithm is purely fictitious and the number of iterations in Equation 14 corresponds to the width of the narrow-band (NB) around the 0-level set where the velocities are initialized. A semi-implicit scheme described in [mikula_inflow-implicitoutflow-explicit_2014, mikula_new_2010] is used here to solve the level set advection equation. Using this method, the equation is written in an alternative form
@@
In order to discretize Equation 16, we use the same notation as in [mikula_inflow-implicitoutflow-explicit_2014]. Consider $\mathrm{p}$ to be a finite volume of a cell and $e_{\mathrm{p}\mathrm{q}}$ the edge between $\mathrm{p}$ and $\mathrm{q}$, $\mathrm{q}\in\mathcal{N}(\mathrm{p})$ where $\mathcal{N}(\mathrm{p})$ is the set of neighboring finite volumes. The length of $e_{\mathrm{p}\mathrm{q}}$ is normalized to $1$. Let $\mathbf{n}_{\mathrm{p}\mathrm{q}}$ be the outer normal vector to $e_{\mathrm{p}\mathrm{q}}$ with respect to $\mathrm{p}$. Finally, let us denote $\bar{\phi}_{\mathrm{p}}$ the constant reconstruction of $\phi_{\mathrm{p}}$ in the finite volume $p$ and $\bar{\phi}_{\mathrm{p}\mathrm{q}}$ the constant reconstruction of $\phi$ on $e_{\mathrm{p}\mathrm{q}}$. Integrating volume $\mathrm{p}$, Equation 16 yields
@@
where $\tau$ is the time step, $\Delta$ the uniform grid spacing and $n$ a given time step. To reconstruct of $\bar{\phi}_{\mathrm{p}}$, $\bar{\phi}_{\mathrm{p}\mathrm{q}}$ and $|\nabla\phi_{\mathrm{p}\mathrm{q}}|$, we use the diamond-cell strategy described in [mikula_new_2010]. This evolution equation however does not preserve the signed distance property, which can be restored by periodically employing a reinitialization procedure by solving an Eikonal equation. There exist many numerical methods for solving this equation and, here we have adopted that of Min [min_reinitializing_2010], also recently used in the work of Limare [Limare2022]. For further details on the numerical implementation of the level set coupled to a cut cell method, we refer the reader to [fullana2022, Fullana2023, QuirosRodriguez2024].
@@ 3 Adjoint problem
We will now proceed with the derivation of the corresponding continuous adjoint problem. The main novelty behind our derivation is that the velocity field $\mathbf{u}$, known at each time step from the forward solution, is treated as a known variable in the adjoint equations. This allows us to discard the adjoint of the Navier–Stokes equations in the fluid phase and leads to an incomplete adjoint formulation. The minimization procedure will serve as a validation of this hypothesis.
@@
Turning now to the adjoint level set problem; due to the geometric non-linearity induced by $\phi$ on the domains of integration a careful treatment of each term is required as described in [bernauer_optimal_2011]. Denoting $D\llbracket H(\phi);\delta\phi\rrbracket$ the variation of $H(\cdot)$ in the direction $\delta\phi$, the adjoint level set equations are derived by setting $\mathcal{L}_{\phi}(\cdot)=0$
@@
This equation can be interpreted as a first-order conservation law on $\Gamma$ and the source term on the right-hand-side can be extended in the neighborhood of $\Gamma$. For details on the derivation and simplifications of $\mathcal{L}_{\phi}(\cdot)=0$, we refer the reader to [fullana2022].
@@ 4 Optimization results
Adjoint-based methods—being efficient gradient-based optimization approaches—are often the methods of choice in PDE-constrained optimal control. By solving an adjoint equation once, one obtains the gradient of the objective with respect to all control variables at minimal cost, making the optimal-control formulation both tractable and efficient even for large-scale parameter spaces. We choose to solve the minimization problem (MP) by using the limited memory BFGS (L-BFGS) method, a quasi-Newton method originally described in [Liu1989]. The main characteristic of this method is that it determines the descent direction by preconditioning the gradient with an approximation of the Hessian matrix. This information is obtained using past approximations – the number of approximations is determined by the memory length parameter which is set to $m=10$ – as well as the gradient. As an initial guess for the initial Hessian, we use the scaled identity matrix as described in [Wright2006]. Algorithm 1 summarizes the L-BFGS method used in our numerical example. The algorithm is stopped at a given iteration $n$ if one of the following criteria is fulfilled
@@
  1exSolve the forward Stefan problem (2) for $T^{k}$ and $\phi^{k}$
  1exSolve the adjoint Stefan problem (3) for $\Theta^{k}$ and $\psi^{k}$
  1exCompute the gradient:
@@
  1exChoose an initial approximation to the inverse of the Hessian $H^{k}_{0}$
  1exConstruct the direction $d^{k}=-H^{k}\nabla\mathcal{J}^{k}$
  1exDetermine $\sigma^{k}$ using a Line Search algorithm with backtracking where $\sigma^{k}=\operatorname{argmin}\mathcal{J}(w^{k}+\sigma^{k}d^{k})$
  1exUpdate $w^{k+1}=w^{k}+\sigma^{k}d^{k}$
[…]
@@
We now compare the results to a derivative-free method (where no information on the gradient is required), the Particle Swarm optimization algorithm [ParticleSwarm]. This method attempts to improve global convergence by switching between four evolutionary states: exploration, exploitation, convergence, and jumping out. In the jumping out state it tries to take the best particle and move it away from its local optimum, to improve the ability to find a global one. We test the derivative-free optimization procedure on the cases presented using the same parameters for the cost functional. Table 1 summarizes the comparison between both methods. As expected, the derivative-free one attains a lower minimum at a cost of a much higher number of function evaluations (one order of magnitude higher). The L-BFGS algorithm driven by the incomplete adjoint thus attains objective values close to those of the particle-swarm optimizer at a fraction of the computational cost.